OpenStudy (anonymous):

What is the equation written in slope intercept form? 9x-6y=102 ?

OpenStudy (anonymous):

9x-6y=102 6y=9x-102 y=(3/2)x-17

OpenStudy (anonymous):

Ok I see thanks!